OpenStudy (anonymous):

what is 6*10^24 --------- 8*10^8

OpenStudy (jdoe0001):

\(\bf \cfrac{6\times 10^{24}}{8\times 10^8}\implies \cfrac{\cancel{ 6 }}{\cancel{ 8 }}\times \cfrac{10^{\cancel{ 24 }}}{10^{\cancel{ 8 }}}\implies ?\)

OpenStudy (anonymous):

help